Checl Suppose that Casino Royale has issued bonds that mature in 1 year. They currently offer a yield of 29%. However, there is a 50% chance that Casino will default and bondholders will receive nothing. What is the expected yield on the bonds? Assume these are zero coupon bonds with annual compounding. (Input the amount as a positive value and as a percent rounded to 1 decimal place.) of < Prev 100123 喆 Ney, >
Answer #1

Par value = 1000
Price of bond = 1000/(1+29%) =775.19

Expected Value of Bond = 50%*1000 +50%*0 = 500

The Expected yield = (500-775.19)/775.19 = -35.50%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well